# Predicting Networks Before They Happen: Experimentation on a Real-Time V2X Digital Twin

### v1

@@ I Introduction
Connected and automated vehicles are pushing wireless networks to operate increasingly close to their physical limits, both in terms of reliability and reaction time. Safety critical Vehicle-to-Everything (V2X) applications such as collision avoidance, cooperative perception, and automated intersection management require the network to reason not only on the current state of the environment, but also on how positions and channels will evolve in the immediate future, from tens of milliseconds to a few seconds ahead [11038674]. This is precisely the promise of Network Digital Twins (NDTs): a software replica of the physical system, continuously synchronized with live data, that can be queried to evaluate “what–if” situations before they affect real users [9839640]. Despite significant progress, most existing NDTs operate in an essentially offline fashion. High–fidelity ray tracing, realistic mobility engines, and multi-RAT protocol stacks are routinely used to obtain accurate performance evaluations, but without strict timing constraints, simulation runs may take seconds, minutes, or even hours to complete. This is acceptable for design-time analysis, but incompatible with a reactive or predictive Digital Twin (DT), where results must be delivered within the tight latency budget [9429703] imposed by V2X applications, as an example. If the twin answers too late, vehicles will have moved, the channel will have changed, and any recommended action may already be obsolete or even unsafe. Two recent platforms partially address this gap. The Tokyo Mobility DT—from here on referred to as Mobility DT—developed by the Institute of Science Tokyo, Japan [10443037] provides a detailed 3D reconstruction of the Ookayama Campus area in Tokyo, Japan supporting real or replayed trajectories and enabling realistic urban-scale mobility scenarios. Conversely, VaN3Twin [pegurri2025van3twinmultitechnologyv2xdigital] extends the ms-van3t/ns-3 framework [ms-van3t-journal-2024] by integrating Sionna RT [sionna] in-the-loop, thereby enabling deterministic ray-traced channel prediction and the implementation of Intelligent Transportation Systems (ITS) protocols over full-stack realizations of multiple Radio Access Technologies (RATs). The resulting framework has been validated against on-field measurements, demonstrating close agreement between predicted and observed channel behavior [pegurri2025van3twinmultitechnologyv2xdigital]. Each platform solves a different part of the problem: the Mobility DT offers realism in geometry and mobility, while VaN3Twin offers realism in the complete communication modeling. Neither, however, provides an end-to-end and real-time pipeline capable of ingesting live mobility, predicting near-future trajectories, computing near-future channels, and request action within a strict deadline.
@@ I-A State of the art
General-purpose network simulators such as ns-3 and OMNeT+ form the foundation of most V2X simulation frameworks. Among them, ms-van3t is one of the most comprehensive: it supports IEEE 802.11p, LTE-V2X, NR-V2X, and the full ETSI ITS-G5 facilities layer, and integrates with SUMO and CARLA for mobility simulation [ms-van3t-journal-2024]. Several recent works interpret such simulators as components of broader vehicular NDTs. Examples include blockage-aware mmWave DTs for multi-hop topologies [roongpraiwan2025digital] and multi-modal DTs combining mobility, sensing, and coarse communication abstractions [CazzellaV2XDT]. These platforms move beyond offline simulation but still rely on simplified PHY abstractions to maintain responsiveness. Traditional wireless propagation models in ns-3 and OMNeT++ are stochastic (e.g., 3GPP TR 36.885/38.901) or based on analytical families such as WINNER, able to capture large-scale trends but not the site-specific multipath structure. This limitation becomes critical in DT settings, where small changes in geometry or mobility strongly influence short-term channel evolution. To overcome this, several works integrate external ray-based channel simulators into network simulators. MATLAB-based approaches [matlab-rt-colosseum] emphasize PHY accuracy and hardware emulation, but lack full V2X stacks and scale poorly. Other works import offline ray-traced maps into ns-3 [gaugel2012accurate], enabling bit-level simulation but preventing synchronized updates. OPAL–OMNeT++ integration [RUZNIETO2023100964] achieves realistic geometry but targets non-V2X technologies. VaN3Twin, however, advances this direction by coupling ms-van3t with Sionna RT into a comprehensive V2X NDT framework, enabling high-fidelity, in-the-loop ray-traced propagation and realistic modeling of multi-RAT coexistence over shared spectrum portions. However, despite its channel realism, VaN3Twin has primarily been used offline or in emulation mode, without enforcing a strict real-time prediction deadline. GPU-accelerated ray tracing software such as Sionna RT now enable repeated evaluations with competitive runtime [zhu2024toward]. This has motivated new in-the-loop architectures, including ns3sionna [ns3-sionna-falko] for Wi-Fi and our prior ns-3 + Sionna RT integration [Pegu2505:Toward] for multi-RAT 6G studies. DT-CoVeSS [Twardokus2505:DT-CoVeSS] similarly uses Sionna RT for high-fidelity analysis of C-V2X security. These works demonstrate that in-the-loop ray tracing integration is feasible, but none targets city-scale mobility, full V2X communication stacks, and strict prediction deadlines simultaneously. Vehicular DTs outside ns-3 focus increasingly on near real-time operation. RAVEN [RAVEN-paper] combines live data and contextual information to predict mmWave conditions and support interference management. Vehicle-to-Cloud ADAS [V2C-ADAS] demonstrates cloud-assisted actuation using DT logic. Large-scale mobility DTs such as TuST [TuST-paper-journal] and indoor crowd models [9861008] show that realistic spatio-temporal digital replicas can be built from real data. However, these systems often rely on simplified communication models or separate mobility and communication loops. What is missing is a DT that combines: (i) realistic, city-scale mobility; (ii) a full-stack, ray-traced V2X simulator; and (iii) an explicit timing budget ensuring that predictions are delivered before the corresponding physical time occurs.
@@
Paper Contributions To address the gap identified above, we integrate VaN3Twin within the Tokyo Mobility DT ecosystem using Sionna RT as a shared ray tracing core, orchestrating data exchange and scheduling to enable short-term future channel prediction at latencies compatible with real-time V2X operation. Predictions that arrive too late are discarded to ensure consistency with the evolving physical system. This integration exposes the key challenges of real-time NDT operation, including the computational cost of ray-traced propagation, the synchronization of mobility and geometry between platforms, and the scheduling required to deliver actionable predictions before the corresponding physical time occurs. Rather than proposing new trajectory or channel prediction algorithms, our goal is to quantify how far a real system can be pushed toward real-time predictive operation when combining a live mobility testbed with a high-fidelity V2X DT.
@@
  • We introduce a timing-constrained workflow for future channel state prediction, where mobility updates, 3D scene synchronization, ray tracing, and network simulation are orchestrated to meet a real-time budget;
@@ II Integration architecture
This section presents the architecture of the proposed framework and details the interactions among its main components. As shown in Figure 1, the framework integrates three core components: (i) The Physical Entities, i.e., the vehicles equipped with sensing and communication capabilities; (ii) The Mobility DT, which maintains a live digital replica of the surrounding traffic and predicts short-term future mobility states; (iii) VaN3Twin, the NDT engine responsible for full-stack replication and prediction of V2X communication processes. Physical Entities provide time-stamped measurements to the Mobility DT, which produces short-term mobility predictions for VaN3Twin. VaN3Twin uses these predictions to simulate future communication states and returns per-link KPIs. In this architecture, the Mobility DT serves as the coordinating component of the system. It aggregates real-time mobility and perception data collected by LiDAR-equipped Physical Entities as well as by LiDAR and camera sensors mounted on nearby Road Side Units (RSUs) or, eventually, network-enabled positioning techniques [10644093]. Using this information, it continuously maintains and forecasts the mobility state of all dynamic elements in the scenario. Mobility updates are processed at a fixed, pre-defined rate, and predictions are generated for a single, short-term horizon $h$, selected to remain compatible with the end-to-end latency budget of the system. The prediction horizon is kept fixed during operation to ensure deterministic scheduling and bounded computation time. The Mobility DT then orchestrates network-level prediction requests to VaN3Twin and subsequently disseminates predicted communication states back to the Physical Entities—possibly including control commands to be applied to the network configuration. Communication between system elements is structured according to their functional roles. A specific physical entity provides the interface between Physical Entities and the Mobility DT, leveraging a 2.4 GHz IEEE 802.11n Wi-Fi Control Plane (CP) for low latency over-the-air control operations. Meanwhile, the Mobility DT and VaN3Twin communicate over a wired Ethernet Local Area Network (LAN), reflecting their co-location at the network edge to minimize latency. All inter-component exchanges employ lightweight UDP messages with JSON-formatted payloads, enabling efficient and low-overhead transmission of state and control information. Message delivery is best-effort: occasional packet losses are tolerated and naturally overridden by subsequent updates, avoiding retransmission delays. All components rely on a shared time base, with timestamps carried in each message and clocks synchronized using standard network time synchronization mechanisms, ensuring consistent temporal alignment across mobility and communication predictions. The detailed data exchange loop is described in the following subsection.
@@ II-A End-to-end Workflow
The end-to-end workflow illustrated in Figure 2 specifies all the interactions introduced above and clarifies how mobility prediction and communication prediction are coupled in practice. Physical Entities periodically transmit their sensed location and measured communication data to the Mobility DT—located in the edge cloud—with an implementation-dependent reporting period $\Delta t_{\mathrm{PE}}$. These updates are time-stamped and provide the input required to keep the digital replica synchronized with the physical system and to continuously adjust the internal prediction models. Based on the aggregated and synchronized mobility state, the Mobility DT computes short-term trajectory forecasts for all relevant dynamic objects. Network prediction requests toward VaN3Twin are then generated according to a configurable policy, which may be periodic or event-driven, and always target a single future time instant $t+h$. The prediction horizon $h$ is chosen so as to remain compatible with the end-to-end latency budget; multi-horizon requests are also possible in principle but are outside the scope of this work. Upon receiving a request, VaN3Twin instantiates a future communication scenario corresponding to the target time $t+h$. Predicted vehicle poses are used to construct a consistent 3D scene, which serves as input to the deterministic ray-based channel model formalized in Section III. The resulting channel realizations drive full-stack V2X simulations, possibly yielding protocol- and packet-level performance indicators. In parallel, the recently observed communication measurements are used to compensate for systematic prediction biases, such as offsets on the computed received power. The predicted communication state is then returned to the Mobility DT and made available for network-level decision making. When required, the DT translates these predictions into control actions which are conveyed back to the Physical Entities over the dedicated Wi-Fi CP, thereby closing the loop between the digital and physical domains. To reason about the limits of this approach, the next section provides a formal characterization of the proposed integration, detailing the underlying trajectory prediction model, the channel estimation process, and the resulting trade-offs between model complexity, end-to-end latency, and the maximum achievable prediction horizon.
@@ III End-to-End Analysis
This section provides a formal characterization of the integration presented in Section II. We describe the trajectory prediction model, formalize the ray-based channel prediction process, and introduce a compact notion of model complexity that links physical-model fidelity to end-to-end latency and feasible prediction horizons.
@@ III-A Mobility and Channel Prediction
Each vehicle $i$ in the scenario is represented by the kinematic state $\mathbf{x}^{i}_{t}=(\mathbf{p}^{i}_{t},\,\mathbf{v}^{i}_{t},\,\psi^{i}_{t})$. The state evolution assumes a constant-velocity motion model, with $\mathbf{F}^{i}_{t}$ defined accordingly based on the sampling interval. $\mathbf{z}^{i}_{t}$ represents the state measurement vector, which the DT fuses into a unified state observation. The temporal evolution of each vehicle follows a linear state-space model defined by the state-transition matrix $\mathbf{F}^{i}_{t}$, the control-input matrix $\mathbf{B}^{i}_{t}$, and the measurement model matrix $\mathbf{M}^{i}_{t}$, with process noise covariance $\mathbf{Q}^{i}_{t}$ and measurement noise covariance $\mathbf{R}^{i}_{t}$. The process noise covariance $\mathbf{Q}^{i}_{t}$ accounts for deviations from ideal constant-velocity motion and is treated as a design parameter, whereas the measurement noise covariance $\mathbf{R}^{i}_{t}$ is sensor-dependent and reflects the heterogeneous accuracy of the contributing sensing modalities. Given the filtered estimate at time $t$, the DT propagates the state $h$ steps forward to obtain the predicted pose $\hat{\mathbf{x}}^{\,i}_{t+h|t}$, which defines the forecasted traffic geometry used for channel prediction (Algorithm 1).
@@
Once the future scene configuration is assembled, the DT evaluates the wireless channel through a deterministic ray-based propagation model. For each transmitter–receiver pair $(i,j)$, the ray tracing engine emits $N_{\mathrm{rays}}$ rays from the antenna of vehicle $i$, with departure directions covering the unit sphere. Each ray propagates through the environment and may undergo up to $N_{\mathrm{int}}$ interactions with surrounding surfaces. A ray reaching the antenna of vehicle $j$ constitutes a valid propagation path and is added to the set $\mathcal{P}^{i,j}_{\mathrm{valid}}$. Each path $p\in\mathcal{P}^{i,j}_{\mathrm{valid}}$ is characterized by its length $d_{p}^{i,j}$, delay $\tau_{p}^{i,j}=d_{p}^{i,j}/c$, angles of departure $(\theta^{(p)}_{\mathrm{T}},\varphi^{(p)}_{\mathrm{T}})$ and arrival $(\theta^{(p)}_{\mathrm{R}},\varphi^{(p)}_{\mathrm{R}})$, and by an ordered sequence of $N_{p}$ interactions. Its effect on the received field is described by the $2\times 2$ propagation matrix:

### v2

@@ I Introduction
Connected and automated vehicles are pushing wireless networks to operate increasingly close to their physical limits, both in terms of reliability and reaction time. Safety critical Vehicle-to-Everything (V2X) applications such as collision avoidance, cooperative perception, and automated intersection management require the network to reason not only about the current state of the environment, but also on how positions and channels will evolve in the immediate future, from tens of milliseconds to a few seconds ahead [4]. This is precisely the promise of Network Digital Twins (NDTs): a software replica of the physical system, continuously synchronized with live data, that can be queried to evaluate “what–if” situations before they affect real users [6]. Despite significant progress, most existing NDTs operate in an essentially offline manner. High–fidelity ray tracing, realistic mobility engines, and multi-RAT protocol stacks are routinely used to obtain accurate performance evaluations, but without strict timing constraints, simulation runs may take seconds, minutes, or even hours to complete. This is acceptable for design-time analysis, but incompatible with an online, predictive Digital Twin (DT), for which results must be delivered within the tight latency budget [20], for example as imposed by V2X applications. If the twin delivers its prediction too late, vehicles will have moved, the channel will have changed, and any recommended action may already be obsolete or even unsafe. Two recent platforms partially address this gap. The Tokyo Mobility DT—from here on referred to as Mobility DT—developed by the Institute of Science Tokyo, Japan [18] provides a detailed 3D reconstruction of the Ookayama Campus area in Tokyo, Japan. It supports both real-time and replayed trajectories enabling realistic urban-scale mobility scenarios. Conversely, VaN3Twin [9] extends the ms-van3t/ns-3 framework [13] with the in-the-loop integration of Sionna RT [5], thereby enabling deterministic ray-based channel prediction and the implementation of Intelligent Transportation Systems (ITS) protocols over full-stack realizations of multiple Radio Access Technologies (RATs). The resulting framework has been validated against on-field measurements, demonstrating close agreement between predicted and observed channel behavior [9].
Each platform solves a different part of the problem: the Mobility DT offers realism in geometry and mobility, while VaN3Twin offers realism across the complete communication stack. Neither, however, provides an end-to-end real-time pipeline capable of ingesting live mobility, predicting near-future trajectories, computing near-future channels, and take action within a strict deadline.
@@ I-A State of the art
General-purpose network simulators such as ns-3 and OMNeT++ form the foundation of most V2X simulation frameworks. Among them, ms-van3t is one of the most comprehensive: it supports IEEE 802.11p, LTE-V2X, NR-V2X, and the full ETSI ITS-G5 facilities layer, and integrates with SUMO and CARLA for mobility simulation [13]. Several recent works interpret such simulators as components of broader vehicular NDTs. Examples include blockage-aware mmWave DTs for multi-hop topologies [14] and multi-modal DTs combining mobility, sensing, and coarse communication abstractions [1]. These platforms move beyond offline simulation but still rely on simplified PHY abstractions to maintain responsiveness. Traditional wireless propagation models in ns-3 and OMNeT++ are stochastic (e.g., 3GPP TR 36.885/38.901) or based on analytical families such as WINNER, able to capture large-scale trends but not the site-specific multipath structure. This limitation becomes critical in DT settings, where small changes in geometry or mobility strongly influence short-term channel evolution. To overcome this, several works integrate external ray-based channel simulators into network simulators. MATLAB-based approaches [15] emphasize PHY accuracy and hardware emulation, but lack full V2X stacks and exhibit poor scalability. Other works import offline ray-traced maps into ns-3 [3], enabling bit-level simulation but preventing synchronous updates. OPAL–OMNeT++ integration [16] achieves realistic geometry but targets non-V2X technologies. VaN3Twin, however, advances this direction by coupling ms-van3t with Sionna RT into a comprehensive V2X NDT framework, enabling high-fidelity, in-the-loop ray-traced propagation and realistic modeling of multi-RAT coexistence over shared spectrum portions. However, despite its channel realism, VaN3Twin has primarily been used offline or in emulation mode, without enforcing a strict real-time prediction deadline. GPU-accelerated ray tracing software such as Sionna RT now enable repeated evaluations with competitive runtime [21]. This has motivated new in-the-loop architectures, including ns3sionna [22] for Wi-Fi and our prior ns-3 + Sionna RT integration [10] for multi-RAT 6G studies. DT-CoVeSS [17] similarly uses Sionna RT for high-fidelity analysis of CV2X security. These works demonstrate that in-the-loop ray tracing integration is feasible, but none targets city-scale mobility, full V2X communication stacks, and strict prediction deadlines simultaneously. Vehicular DTs outside ns-3 focus increasingly on near real-time operation. RAVEN [2] combines live data and contextual information to predict mmWave conditions and support interference management. Vehicle-to-Cloud ADAS [19] demonstrates cloud-assisted actuation using DT logic. Large-scale mobility DTs such as TuST [11] and indoor crowd models [12] show that realistic spatio-temporal digital replicas can be built from real data. However, these systems often rely on simplified communication models or separate mobility and communication loops. What is missing is a DT that combines: (i) realistic, city-scale mobility; (ii) a full-stack, ray-traced V2X simulator; and (iii) an explicit timing budget ensuring that predictions are delivered before the event takes place in the physical world.
@@
Paper Contributions To address the gap identified above, we integrate VaN3Twin within the Tokyo Mobility DT ecosystem using Sionna RT as a shared ray tracing core, orchestrating data exchange and scheduling to enable short-term future channel prediction at latencies compatible with real-time V2X operation. Any predictions that arrive too late are discarded to maintain consistency with the evolving physical system. This integration exposes the key challenges of real-time NDT operation, including the computational cost of ray-traced propagation, the synchronization of mobility and geometry between platforms, and the scheduling required to deliver actionable predictions before the corresponding physical time occurs. Rather than proposing new trajectory or channel prediction algorithms, our goal is to quantify how far a real system can be pushed toward real-time predictive operation when combining a live mobility testbed with a high-fidelity V2X DT.
@@
  • We introduce a timing-constrained workflow for future channel state prediction, where mobility updates, 3D scene synchronization, ray tracing, and network simulation are orchestrated to meet a strict latency budget;
@@ II Integration architecture
This section presents the architecture of the proposed framework and details the interactions among its main components. As shown in Figure 1, the framework integrates three core components: (i) The physical entities, i.e., the vehicles equipped with sensing and communication capabilities; (ii) The Mobility DT, which maintains a live digital replica of the surrounding traffic and predicts short-term future mobility states; (iii) VaN3Twin, the NDT engine responsible for full-stack replication and prediction of V2X communication processes. Physical entities provide time-stamped measurements to the Mobility DT, which produces short-term mobility predictions for VaN3Twin. VaN3Twin uses these predictions to simulate future communication states and returns per-link KPIs. In this architecture, the Mobility DT serves as the coordinating component of the system. It aggregates real-time mobility and perception data collected by LiDAR-equipped physical entities as well as by LiDAR and camera sensors mounted on nearby Road Side Units (RSUs) or, eventually, network-enabled positioning techniques [7]. Using this information, it continuously maintains and forecasts the mobility state of all dynamic elements in the scenario. Mobility updates are processed at a fixed, pre-defined rate, and predictions are generated for a single, short-term horizon $h$, selected to remain compatible with the end-to-end latency budget of the system. The prediction horizon is kept fixed during operation to ensure deterministic scheduling and bounded computation time. The Mobility DT then orchestrates network-level prediction requests to VaN3Twin and subsequently disseminates predicted communication states back to the physical entities—possibly including control commands to be applied to the network configuration. Communication between system elements is structured according to their functional roles. A specific physical entity provides the interface between physical entities and the Mobility DT, leveraging a 2.4 GHz IEEE 802.11n Wi-Fi Control Plane (CP) for low-latency over-the-air control operations. Meanwhile, the Mobility DT and VaN3Twin communicate over a wired Ethernet Local Area Network (LAN), reflecting their co-location at the network edge to minimize latency. All inter-component exchanges employ lightweight UDP messages with JSON-formatted payloads, enabling efficient and low-overhead transmission of state and control information. Message delivery is best-effort: occasional packet losses are tolerated and naturally compensated for by subsequent updates, avoiding retransmission delays. All components rely on a shared time base, with timestamps carried in each message and clocks synchronized using standard network time synchronization mechanisms, ensuring consistent temporal alignment across mobility and communication predictions. The detailed data exchange loop is described in the following subsection.
@@ II-A End-to-end Workflow
The end-to-end workflow illustrated in Figure 2 specifies all the interactions introduced above and clarifies how mobility prediction and communication prediction are coupled in practice. Physical entities periodically transmit their sensed location and measured communication data to the Mobility DT—located in the edge cloud—with an implementation-dependent reporting period $\Delta t_{\mathrm{PE}}$. These updates are time-stamped and provide the input required to keep the digital replica synchronized with the physical system and to continuously adjust the internal prediction models. Based on the aggregated and synchronized mobility state, the Mobility DT computes short-term trajectory forecasts for all relevant dynamic objects. Network prediction requests to VaN3Twin are then generated according to a configurable policy, which may be periodic or event-driven, and always target a single future time instant $t+h$. The prediction horizon $h$ is chosen so as to remain compatible with the end-to-end latency budget; multi-horizon requests are also possible in principle but are outside the scope of this work. Upon receiving a request, VaN3Twin generates a future communication scenario corresponding to the target time $t+h$. Predicted vehicle poses are used to construct a consistent 3D scene, which serves as input to the deterministic ray-based channel model formalized in Section III. The resulting channel realizations drive full-stack V2X simulations, possibly yielding protocol- and packet-level performance indicators. In parallel, the recently observed communication measurements are used to compensate for systematic prediction biases, such as offsets in the computed received power. The predicted communication state is then returned to the Mobility DT and made available for network-level decision making. When required, the DT translates these predictions into control actions which are conveyed back to the physical entities over the dedicated Wi-Fi CP, thereby closing the loop between the digital and physical domains. To reason about the limits of this approach, the next section provides a formal characterization of the proposed integration, detailing the underlying trajectory prediction model, the channel estimation process, and the resulting trade-offs between model complexity, end-to-end latency, and the maximum achievable prediction horizon.
@@ III End-to-End Analysis
This section provides a formal characterization of the integration presented in Section II. We describe the trajectory prediction model, formalize the ray-based channel prediction process, and introduce a concise characterization of model complexity that links physical-model fidelity to end-to-end latency and feasible prediction horizons.
@@ III-A Mobility and Channel Prediction
Each vehicle $i$ in the scenario is represented by the kinematic state $\mathbf{x}^{i}_{t}=(\mathbf{p}^{i}_{t},\,\mathbf{v}^{i}_{t},\,\psi^{i}_{t})$. The state evolution assumes a constant-velocity motion model, with $\mathbf{F}^{i}_{t}$ defined accordingly based on the sampling interval. $\mathbf{z}^{i}_{t}$ represents the state measurement vector, which the DT fuses into a unified state estimate. The temporal evolution of each vehicle follows a linear state-space model defined by the state-transition matrix $\mathbf{F}^{i}_{t}$, the control-input matrix $\mathbf{B}^{i}_{t}$, and the measurement model matrix $\mathbf{M}^{i}_{t}$, with process noise covariance $\mathbf{Q}^{i}_{t}$ and measurement noise covariance $\mathbf{R}^{i}_{t}$. The process noise covariance $\mathbf{Q}^{i}_{t}$ accounts for deviations from ideal constant-velocity motion and is treated as a design parameter, whereas the measurement noise covariance $\mathbf{R}^{i}_{t}$ is sensor-dependent and reflects the heterogeneous accuracy of the contributing sensing modalities. Given the filtered estimate at time $t$, the DT propagates the state $h$ steps forward to obtain the predicted pose $\hat{\mathbf{x}}^{\,i}_{t+h|t}$, which defines the forecasted traffic geometry used for channel prediction (Algorithm 1).
@@
Once the future scene configuration is assembled, the DT evaluates the wireless channel through a deterministic ray-based propagation model. For each transmitter–receiver pair $(i,j)$, the ray tracing engine emits $N_{\mathrm{rays}}$ rays from the antenna of vehicle $i$, with departure directions covering the unit sphere. Each ray propagates through the environment and may undergo up to $N_{\mathrm{int}}$ interactions with surrounding surfaces. A ray reaching the antenna of vehicle $j$ is part of a valid propagation path that is added to the set $\mathcal{P}^{i,j}_{\mathrm{valid}}$. Each path $p\in\mathcal{P}^{i,j}_{\mathrm{valid}}$ is characterized by its length $d_{p}^{i,j}$, delay $\tau_{p}^{i,j}=d_{p}^{i,j}/c$, angles of departure $(\theta^{(p)}_{\mathrm{T}},\varphi^{(p)}_{\mathrm{T}})$ and arrival $(\theta^{(p)}_{\mathrm{R}},\varphi^{(p)}_{\mathrm{R}})$, and by an ordered sequence of $N_{p}$ interactions. Its effect on the received field is described by the $2\times 2$ propagation matrix: